Solve for x e^-2+e^2+x=2
Answer by RAY100(1637)

e^-2 +e^2 +x=2
e^2 = 7.389 (calculator)
1/(e^2) +e^2 +x=2
(1/7.389)+7.389+x=2
subt2 and x both sides
(1/7.389) +7.389 -2 =(-x)
.135 +7.389 -2 = (-x)
5.514 = (-x)
mult by (-1) both sides
(-5.514) = x answer
check
e^-2 +e^2 +x=2
.135 + 7.389 +(-5.514) =2
2=2 ok